Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370


News, Travel, Visitors (McMullin)

  • Overall status: In B configuration; 4-band dipoles are coming down tomorrow. Some issues with final antenna positions (baseline determination have been confusing for some antennas).
  • Meerkat staff visit status (Butler) - M&C group included their OPT, online system and correlator software; in the design stage.
  • Key dates:
    • 09-23 Feb: 4-band test installation
    • 16 Feb: Correlator Fire Suppression tests - Complete; worked 2 out of 3(!).
    • 07-10 Mar: New Worlds, New Horizons (Santa Fe)
    • 07 May - 15 Jun: 4-band science observing
  • Staff Travel:
  • Visitors:
    • 05 Jan - 06 Apr: John Cannon sabbatical at DSOC
    • 22 Feb - 08 Mar: Sonja Vrcic (Correlator)
    • 28 Feb - 28 Mar: Bob Sault

Commissioning priorities for the week (McMullin, Chandler)

Double maintenance day week:
  • Mon
    • All day: Science Operations/testing observations
  • Tue
    • 0900-1700: Testing/software development time
    • 1800-0630: Science operations
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations
  • Thu, Fri:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations
  • Sat-Sun
    • Science operations + planned testing

Issues for Science Observing:
  • Correlator: fshift sub-band anti-aliasing causing phantom spectral lines when using recirculation for narrow bands (only worry if Michael has to prep your file)
  • Correlator: Missing/mislabeled scans in data sets (small percentage but recurring).
  • Correlator: time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags);fix available pending testing verification.
  • Construction: RFI sweep at Ku, K, Ka, Q pre-requisite for high pass filters for use with L301.
  • Calibration: Switched power heuristics for reduction.
  • Array/Construction: Dhawan/Rupen: RF switch issues; new set coming to be tested.
  • Correlator: Delay rocking across bands.
  • Correlator: Subset of BlBs misbehaving (only affects hardware stacking modes)
  • Antenna/High Freq: Reference pointing accuracy (C, X, Ku). Rick has noted that we are worse than before - larger fraction of observations indicate larger pointing offsets (Perley).
  • Array: Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
  • WIDAR (Rupen):
    • Notes: Update from last week: still an unresolved issue (BlB did not respond to a configuration from the CM, beginning with the RXP) - roll back confusion caused much of the rest. Some data issues over the weekend (one deformatter lost lock; one station board went bad for a period of time - being investigated today).
  • Computing (Robnett):

Software status and updates (Butler)

Commissioning Reports

Pipeline: Sjouwerman: Pipeline development status and plans

Antenna: Jackson: ACU Meeting summary

  • Working on 1 month timescale to prepare a report for a new drive system and deploy to one antenna, includes modifications to ACU, FRM and FRM motors. Have available pending funding but will be able to do array in a piecemeal fashion. This won't provide faster slew times but will enable faster settling so more time on source through observing (current number is about 6 seconds to settle can get down to 3-4 seconds).

Group updates (Low frequency, Spectral line, 4-Band wrap-up, Phased-array wrap-up)

Commissioning Group updates and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!


RSRO B configuration status

Individual reports (updates)


Demo Science Status

W51 Status report: Golap



Misc Actions

  • [Frail/Butler]: Enable restrictions based on proximity to sun; review OST weather trending.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised. - Sault visit * Low Priority [TBD]: Phase (and amp) wobbles; long history.fshift investigation.
-- JosephMcMullin - 2011-02-21
Topic revision: r3 - 2011-03-01, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback